What is the pH of a 0.231 M solution of hydrazine, N2H4? The pKb of hydrazine is 6.02.
Answer #1

Given

Hydrazine concentration = [N2H4] = 0.231 M

pKb = 6.02

We know that

pKb = -log(Kb)

hence

Kb = 10-pKb = 10-6.02 = 9.55 \times 10-7

Now lets see the dissocition nreaction of hydrazine in water

N2H4 + H2O \rightleftharpoons N2H5+ + OH-  

lets see the ICE table for above reaction

......................N2H4........+....H2O...... \rightleftharpoons.....N2H5+ ....+.....OH-  

Initial.............0.231 M.................................0.....................0.

Change.............-x...................................... +x..................+x

Equilibrium....0.231-x ....................................x...................x

the equilibrium constant for above reaction is

Kb = [N2H5+][OH-] / [N2H4]

9.55 \times 10-7 = (x)(x) /(0.231-x) = x2 / (0.231-x)

As x is very small than 0.231 , 0.231-x = 0.231

then

9.55 \times 10-7 = x2 / 0.231

x2 = 0.231 \times 9.55 \times 10-7 = 2.20 \times 10-7

by taking the square roots on both sides we get

x = 0.000469 M

hence

[OH-] = 0.000469 M

Now we know that

pOH = -log[OH-]

= -log(0.000469)

= 3.33

Now

pH + pOH = 14

the

pH = 14 - pOH = 14 - 3.33 = 10.67

so our answer is

pH = 10.67